OpenStudy (nightmarenight):

How many solutions are there to the following system of equations? 3x + 7y = –10 15x + 35y = –45

OpenStudy (raden):

what do you think about : 3/15 .... 7/35 ..... -10/-45 ?

OpenStudy (anonymous):

the two lines are parallel but there are no solutions because the coefficients don't match up.

OpenStudy (anonymous):

so the lines are parallel, but distinct

OpenStudy (nightmarenight):

So it's 0? no solution isnt an option...

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

well, no solution would be a better answer, but if no solution is not an option, then i might go with 0

OpenStudy (anonymous):

Use either equation to solve for one variable in terms of another. I will use the first equation to solve for x in terms of y. \[3x+7y=-10\]\[3x=-7y-10\]\[x=\frac{-7}{3}y-\frac{10}{3}\]Now replace x in the second equation with this new expression.\[15x+35y=-45\]\[15(-\frac{7}{3}y-\frac{10}{3})+35y=-45\]\[-35y-50+35y=-45\]The -35y and 35y cancel out so you get:\[-50=-45\]When is -50 equal to -45? Never. So there is no solution. (zero solutions)